1. A sample of 2-butanol has a specific rotation of +3.25. Determine the % ee and the molecular composition of this sample.
Answer #1

1.

Given, Observe specific rotation of 2-butanol = +3.25 degree

Pure specific rotation of 2-butanol = +13 degree

Now,

  • % of enantiomeric excess(ee) = Observe specific rotation / Pure specific rotation * 100

= +3.25 / +13 * 100

=(+) 0.25 * 100

= (+) 25%

Hence, ee = 25% and it is excess of (+) isomer

  • ee 25% means (+) and (-) isomers = (100 - 25)% = 75% ( it is racemic mixture of (+) and (-) isomers)
  • (-) isomer = 75/2 % = 37.5%
  • (+) isomer = (25 + 75/2)% = 52.5%

Because ee is added to (+) isomer as it is excess of (+) isomer.

2.

Given,

Observe specific rotation of 2-butanol = -9.75 degree

Pure specific rotation of 2-butanol = + 13 degree

Now,

  • % enantiomeric excess(ee) = - 9.75/+13 * 100

= (-) 0.75 *100

= (-) 75%

Hence, ee = 75% and it is excess of (-) isomer

Now,

  • ee = 75% means (+) and (-) isomers = (100 - 75)% = 25% (it is racemic mixture of (+) and (-) isomers)
  • (+) isomer = 25/2 % =12.5%
  • (-) isomer = (75 + 25/2)% = 87.5%

Here ee = 75% is added to (-) isomer as it is excess of (-) isomer.

Note:Try to understand it ,if you can understand it then ,hope you will solve any questions like these

And if something confusion,then ask me again in comment box. I am always here to help you.